OpenStudy (anonymous):

How would you solve s/3s+2 + s+3/2s-4 = -2s/3s^2-4s-4 ? If the answer is -6/5 , -1 ?

OpenStudy (anonymous):

is this correct \[\frac{ 5 }{3s}+2+5+\frac{ 3 }{2s }-4=\frac{ -2 }{ 3s^2 }-4s-4\]

OpenStudy (anonymous):

no. \[\frac{ S }{ 3S+2 }+ \frac{ S+3 }{ 2S-4 } = \frac{ -2S }{ 3S ^{2}-4S-4 }\]

OpenStudy (anonymous):

well that 3S^2-4S-4 can be factored

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

yes just one moment

OpenStudy (anonymous):

 I am not sure how you like to factor trinomials